OpenStudy (anonymous):

Tickets for a school play cost $5 if you buy them early, and $7 if you buy them the day of the show. If 200 tickets are sold, and the total amount of money from ticket sales was $1340, how many tickets were purchased early?

OpenStudy (anonymous):

system schmistem lets call \(x\) the number of early tickets, so \(200-x\) is the number of late tickets early tickets brought in \(5x\) dollars and late tickets bring \(7(200-x)\)

OpenStudy (anonymous):

you could try an equation like x+y=200, and 5x+7y=1340

OpenStudy (anonymous):

since the total is \(1340\) you can set \[5x+7(200-x)=1340\] and solve for \(x\)

OpenStudy (anonymous):

thanks satellite can i ask you another question and you answer it please

OpenStudy (anonymous):

sure